OpenStudy (anonymous):

What is the area of a circle with a diameter of 6 cm? Use 3.14 for pi. F. 18.84 G. 28.26 H. 76.43 J. 113.04

OpenStudy (anonymous):

\[A_{Circle} = πr^{2}\]\[r = \frac{d}{2}\]\[r = \frac{6}{2} = 3\]\[A_{Circle} = π(3)^{2} = ?\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Just solve this.\[A_{Circle} = π(3)^{2} = 3.14(3)^{2} = 3.14(9) = ?\]